# Data-adaptive structural change-point detection via isolation

### v1

@@ Abstract
In this paper, a new data-adaptive method, called DAIS (Data Adaptive ISolation), is introduced for the estimation of the number and the location of change-points in a given data sequence. The proposed method can detect changes in various different signal structures; we focus on the examples of piecewise-constant and continuous, piecewise-linear signals. We highlight, however, that our algorithm can be extended to other frameworks, such as piecewise-quadratic signals. The data-adaptivity of our methodology lies in the fact that, at each step, and for the data under consideration, we search for the most prominent change-point in a targeted neighborhood of the data sequence that contains this change-point with high probability. Using a suitably chosen contrast function, the change-point will then get detected after being isolated in an interval. The isolation feature enhances estimation accuracy, while the data-adaptive nature of DAIS is advantageous regarding, mainly, computational complexity and accuracy. The simulation results presented indicate that DAIS is at least as accurate as state-of-the-art competitors.
@@ 1 Introduction
Change-point detection, also known as data segmentation, is the problem of finding abrupt changes in data when, at least, one of their properties over time changes. It has attracted a lot of interest over the years, mostly due to its importance in time series analysis and the wide range of applications where change-point detection methods are needed. These include genomics ([7]), neuroscience ([26]), seismic data ([34]), astronomy ([16]) and finance ([10]).
@@
where $X_{t}$ are the observed data and $f_{t}$ is a one-dimensional deterministic signal with structural changes at certain points. This paper is concerned with piecewise-constant and continuous piecewise-linear signals $f_{t}$, meaning that the changes are in the mean or the slope. We highlight that the proposed algorithm can though be extended to other, possibly more complicated, signal structures. Although detecting changes in the mean is a simple case, as noted by [6] more complex change-point problems that allow changes in properties other than the mean, can be reduced to the segmentation problem we are studying. This can be done by applying a suitable transformation to the data that reveals the changes as those in the mean of the transformed data. For example, [11] and [1] provide methods for the detection of multiple change-points in the second-order (i.e. autocovariance and cross-covariance) structure of possibly high dimensional time series, using Haar wavelets as building blocks. The signal is transformed into a piecewise-constant one with identical change-point locations. We want to emphasise that the novelty of our paper does not come from the structure of changes that the method can be applied to. The importance of our work lies in the data-adaptive nature of the proposed algorithm itself.
@@
The various methods available for solving the problem of change-point detection can be mainly split into two categories, according to whether they are optimization-based methods or they use an appropriately chosen contrast function. The former group includes methods that look for the optimal partition of the data, by performing model selection using a penalization function to avoid overfitting. The latter category’s methods do not search for the globally optimal partition of the signal. Instead, the change-point locations are chosen as the most probable location at the given step of the algorithm. Only some of the already existing methods are mentioned in this introduction but comprehensive overviews and more detailed explanations can be found in [40] and [45].
@@
Starting with the optimization-based methods for detecting changes in the mean when $f_{t}$ is piecewise-constant, one of the most common penalty functions is the Schwarz Information criterion ([37]), which was used by [43] for change-point detection, assuming Gaussian random variables $\epsilon_{t}$ in (1). Under the same assumption, [44] studies an estimator based on least squares. Relaxing this assumption and allowing instead for more general distributions belonging to the exponential family, [22] introduces a dynamic program which uses maximum likelihood estimates of the location of the change-points. [33] introduces an AIC-type criterion for change-point models and [13] shows that while an AIC-like information criterion does not give a strongly consistent selection of the optimal number of change-points, a BIC-like information criterion does. [23] employs dynamic programming to guarantee that the exact global optimum, in terms of creating the segments, is found, while [24] proposes an algorithm, called PELT, whose computational cost is linear in the number of observations. [35] introduces the pDPA algorithm, which includes a pruning step towards complexity reduction. [30] combines the ideas from PELT and pDPA leading to two new algorithms, FPOP and SNIP, which have low computational complexity and, therefore, increased speed.
@@
Focusing now on the category where a contrast function is used, for $f_{t}$ being piecewise-constant the relevant function is the absolute value of the CUSUM statistic. A method that has received a lot of attention is the Binary Segmentation algorithm, as introduced in [41]. At each step, the data sequence is split according to the already detected change-points. However, because of checking for a single change-point in intervals that could have any number of change-points, Binary Segmentation has suboptimal accuracy; many methods have been developed with the scope to improve on such drawbacks. One of these methods is proposed by [18], called WBS, which calculates the value of the contrast function on a large number of randomly drawn intervals, which allows detection of change-points in small spacings. [20] proposes WBS2 which draws new intervals each time a detection occurs. A different approach is adopted by [5] with the NOT algorithm. By choosing the narrowest interval that exceeds a pre-defined threshold, the probability that there is exactly one change-point in each interval increases. The ID method in [2] achieves, first, isolation and then detection of each change-point through an idea based on expanding intervals in a sequential way, starting from both the beginning and the end of the data sequence. The isolation of each change-point maximizes the detection power. [26] proposes a seeded binary segmentation algorithm which uses a deterministic construction of search intervals which can be pre-computed. [12], [14] and [27] consider moving sum statistics, while [8], through a ‘bottom-up’ approach, explores reverse segmentation. This involves the creation of a ‘solution path’ by deleting the change-point with the smallest CUSUM value in the segment determined by its closest left and right neighbors, in order to obtain a hierarchy of nested models.
@@
Methods for detecting more general structural changes can again be split in the same two categories. Focusing on optimization-based methods, [4] considers the estimation of linear models based on the least squares principle. [25] introduces a trend filtering (TF) approach to produce piecewise-linear trend estimates using an $\ell_{1}$ penalty. [15] presents a locally dynamic approach, called CPOP, that finds the best continuous piecewise-linear fit to the data using penalized least squares and an $\ell_{0}$ penalty. [17] proposes the MARS method for flexible regression using splines functions with the degree and knot locations determined by the data. Two methods for optimizing the knot locations and smoothing parameters for least-squares or penalized splines are introduced by [39]. [42] introduces the frequentist information criterion for change-point detection which can detect changes in the mean, slope, standard deviation or serial correlation structure of a signal whose noise can be modeled by Gaussian, Wiener, or Ornstein-Uhlenbeck processes. [38] introduces PLANT, a bottom-up type algorithm that finds the points at which there is potential variation of the slope using a likelihood-based approach, and then recursively merges the adjacent segments. Considering algorithms using a suitably chosen contrast function, the algorithms NOT and ID by [5] and [2], respectively, can be applied to piecewise-linear signals.
There are some methods that do not fall in either of the categories as described above. Considering piecewise-constant signals, [28] proposes FDRSeg, which controls the false discovery rate in the sense that the number of falsely detected change-points is bounded linearly by the number of true jumps. [19] achieves a multiscale decomposition of the data with respect to an adaptively chosen unbalanced Haar wavelet basis, using a transformation of the data called TGUH transform. [29] extends TGUH to piecewise-linear signals.
@@
In this project, we propose a data-adaptive change-point detection method, called DAIS (Data Adaptive ISolation), that attempts to isolate the change-points before detection, by also taking into account the potential true locations of the change-points. This means that the algorithm does not start from a random interval, or the beginning/end of the signal, as most algorithms in the literature, but instead starts checking around a point that there is reason to believe it could be a true change-point. The idea behind the data-adaptive nature of the algorithm arises from the fact that in the case that the unobserved true signal $f_{t}$ in (1) is piecewise-constant, taking pairwise differences between consecutive time steps will reveal a constant signal with the value 0, with ‘spikes’ where the change-points occur. This means that the largest ‘spike’, in absolute value, occurs at the location of the change-point with the largest jump in the sequence. Similarly, in the case that $f_{t}$ is continuous and piecewise-linear, differencing the signal twice will again reveal a signal with ‘spikes’ at the locations of the change-points. We try to use this fact in the observed signal, $X_{t}$. In the algorithm, we identify the location of this ‘spike’, which we will be referring to as the ‘largest difference’ for the rest of the paper, and test around it for possible change-points. The novelty of our work lies in the data-adaptive nature of the algorithm as just described.
The DAIS algorithm uses left- and right-expanding intervals around the location of the largest difference found, denoted $d_{s,e}$, in the interval $[s,e]$ in order to identify the potential change-point that might have caused the spike and so must be close to $d_{s,e}$. This is done in a deterministic way around $d_{s,e}$, expanding once either only to the left or only to the right at each step, in an alternate way. Using expanding intervals, we achieve isolation of the change-points, which is desirable as the detection power of the contrast function is maximized in such cases (see Sections 3.1 and 3.2 for the choice of the contrast function). Due to the alternating sides of the expansions, the location of the largest difference, which we have a reason to believe to be also the location of a change-point, is at the midpoint of the interval being checked (or close to it) after an even number of steps has been performed. This increases the power of the contrast function and so gives an advantage to the detection power of the method. Using an expansion parameter $\lambda_{T}$ (more details in Section 4.2) at most $\lceil(e-s+1)/\lambda_{T}\rceil+1$ steps are necessary to check the whole length of the interval around the largest difference. As soon as a change-point is detected, DAIS restarts on two disjoint intervals, one ending at the start-point of the interval where the detection occurred and one starting from the end-point of that same interval. A more detailed explanation of the algorithm can be found in Section 2.
@@
The paper is organized as follows. Section 2 provides details on the methodology of DAIS. We present a simple example followed by a detailed explanation of the algorithm and of the connection between the location of the largest difference and the isolation aspect. In Section 3, we provide theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations as well as some theoretical results concerning the assumption of Section 2.3. Section 4 includes an explanation of how some parameters of DAIS are selected, some modifications to achieve better performance and a note on computational complexity compared with some competitors. In Section 5, we provide results on simulated data and a comparison to state-of-the-art competitors. In Section 6, two examples using real world data are presented. The first one is about crime data reported daily and the second is weekly data on the Euro to British pound exchange rate. Finally, Section 7 concludes the paper with a summary of the most important findings.
@@ 2.1 Simple example
We begin by presenting a toy example of how DAIS works in practice in the case of $f_{t}$ in (1) being a piecewise-constant signal. We have a sequence of length $T=100$, $\sigma=1$ and a change-point at location $r=65$ with jump of magnitude $1.5$. The data are presented in the left plot of Figure 1 with the true underlying signal plotted in red. For the sake of presentation of this example, the expansion parameter is chosen to be $\lambda_{T}=10$. We first calculate the absolute difference of consecutive observations, $X_{t}$, for $t=1,2,\dots,T$. As can be seen by the plot on the right of Figure 1, in this example the largest difference occurs at the location of the change-point ($d_{1,100}=65$). This means that $\textrm{argmax}_{t=1,2,\ldots,99}\lvert X_{t+1}-X_{t}\rvert=65$. As shown in the left plot of Figure 2, DAIS creates right- and left-expanding intervals around this point and so checks, in this order, the intervals $[65,74]$, $[55,74]$ and $[55,84]$, where detection occurs. Note here that the algorithm does not detect the change-point in the first two intervals being checked, most probably either because the magnitude of the jump compared to $\sigma$ is small, or because the number of observations used is not sufficiently large, or a combination of the two. The change-point is identified at location 65 and then the algorithm restarts in the intervals $[1,55]$ and $[84,100]$. As these two intervals contain no change-points in our example, the largest differences are at points where the absolute value of the difference between the noise of two consecutive observations is large. The largest differences are now detected at points $46$ and $85$ in the two intervals, respectively. The algorithm starts checking the interval $[1,55]$, with the expansions being performed in a similar way as to the smaller interval $[84,100]$, on which we focus for simplicity. For the interval $[84,100]$ the left- and right-expanding subintervals checked are: $[85,94]$, $[74,94]$, and $[84,100]$, as shown in the right plot of Figure 2. Since no change-point is detected in either of the intervals, the algorithm terminates.
[…]
[scale=0.7]Tikz_Plots/toy_example
@@
Figure 1: Left plot: The data (in black) and the underlying signal (in red) used for the toy example. The change-point is at location 65. Right plot: The absolute values of consecutive pairwise differences. The maximum, with location $d_{1,100}$, is marked with a red circle.
[…]
[height = 2.3cm, width = 12cm]Tikz_Plots/tikz_3_plots
@@
Figure 2: Example with one change-point at $r=65$ with jump 1.5. Left plot: The largest difference is at $d_{1,100}=65$ and the change-point is detected when the third expanding interval is checked. After detection the next intervals that are checked are $[1,55]$ and $[84,100]$. Right plot: The left- and right-expanding intervals around the location of the largest difference, $d_{74,100}=85$, when checking $[84,100]$. No change-point is detected.

### v2

@@ Abstract
In this paper, a new data-adaptive method, called DAIS (Data Adaptive ISolation), is introduced for the estimation of the number and the location of change-points in a given data sequence. The proposed method can detect changes in various different signal structures; we focus on the examples of piecewise-constant and continuous, piecewise-linear signals. The novelty of the proposed algorithm comes from the data-adaptive nature of the methodology. At each step, and for the data under consideration, we search for the most prominent change-point in a targeted neighborhood of the data sequence that contains this change-point with high probability. Using a suitably chosen contrast function, the change-point will then get detected after being isolated in an interval. The isolation feature enhances estimation accuracy, while the data-adaptive nature of DAIS is advantageous regarding, mainly, computational complexity. The methodology can be applied to both univariate and multivariate signals. The simulation results presented indicate that DAIS is at least as accurate as state-of-the-art competitors and in many cases significantly less computationally expensive.
@@ 1 Introduction
Change-point detection, also known as data segmentation, is the problem of finding abrupt changes in data when, at least, one of their properties over time changes. It has attracted a lot of interest over the years, mostly due to its importance in time series analysis and the wide range of applications where change-point detection methods are needed. These include genomics ([8]), neuroscience ([32]), seismic data ([40]), astronomy ([10]), and finance ([13]).
@@
where $X_{t}$ are the observed data and $f_{t}$ is a one-dimensional deterministic signal with structural changes at certain points. We highlight, however, that the case when $X_{t}$ is multivariate is also discussed in our paper. We focus on the case of piecewise-constant and continuous piecewise-linear signals $f_{t}$, meaning that the changes are in the mean or the slope. However, the proposed algorithm can be extended to other, possibly more complicated, signal structures. Although detecting changes in the mean is a simple case, as noted by [7], more complex change-point problems that allow changes in properties other than the mean, can be reduced to the segmentation problem we are studying. This can be done by applying a suitable transformation to the data that reveals the changes as those in the mean of the transformed data. For example, [14] and [1] provide methods for the detection of multiple change-points in the second-order (i.e. autocovariance and cross-covariance) structure of possibly high dimensional time series, using Haar wavelets as building blocks. The signal is transformed into a piecewise-constant one with change-point locations identical to those in the original signal. We want to emphasise that the novelty of our paper does not come from the structure of changes that the method can be applied to. Upon correct choice of a contrast function, we can cover more scenarios such as the case of piecewise-quadratic signals (see [6] for more details). The importance and novelty of our work lies mainly in the data-adaptive nature of the proposed algorithm itself, which, at each step, starts searching for change-points in the areas that most prominently include one.
@@
The various methods available for solving the problem of change-point detection can be mainly split into two categories, according to whether they are optimization-based methods or they use an appropriately chosen contrast function. The former group includes methods that look for the optimal partition of the data, by performing model selection using a penalization function to avoid overfitting. The latter category’s methods do not search for the globally optimal partition of the signal. Instead, the change-point locations are chosen as the most probable location at each step of the algorithm. Only some of the already existing methods are mentioned in this introduction but comprehensive overviews and more detailed explanations can be found in [47] and [53].
@@
Starting with the optimization-based methods for detecting changes in the mean when $f_{t}$ is piecewise-constant, one of the most common penalty functions is the Schwarz Information Criterion ([44]), which was used by [51] for change-point detection, assuming Gaussian random variables $\epsilon_{t}$ in (1). Under the same assumption, [52] studies an estimator based on least squares. Relaxing this assumption and allowing instead for more general exponential family distributions, [27] introduces a dynamic programming algorithm which uses maximum likelihood estimates of the location of the change-points. [39] introduces an AIC-type criterion for change-point models and [19] shows that while an AIC-like information criterion does not give a strongly consistent selection of the optimal number of change-points, a BIC-like information criterion does. [28] employs dynamic programming to guarantee that the exact global optimum, in terms of creating the segments, is found, while [30] proposes an algorithm, called PELT, whose computational cost is linear in the number of observations. [42] introduces the pDPA algorithm, which includes a pruning step towards complexity reduction. [36] combines the ideas from PELT and pDPA leading to two new algorithms, FPOP and SNIP, which have low computational complexity.
@@
Focusing now on the category where a contrast function is used, for $f_{t}$ being piecewise-constant the relevant function is the absolute value of the CUSUM statistic, which is defined in Section 3.1. A method that has received a lot of attention is the Binary Segmentation algorithm, as introduced in [48]. It starts by searching the whole data sequence for one change-point. Subsequently, at each step, the data sequence is split according to the already detected change-points. However, because of checking for a single change-point in intervals that could have more than one change-points, Binary Segmentation has suboptimal accuracy; many methods have been developed with the scope to improve on such drawbacks. One of these methods is proposed by [23], called WBS, which calculates the value of the contrast function on a large number of randomly drawn intervals, which allows detection of change-points in small spacings. [25] proposes WBS2 which draws new intervals each time a detection occurs. A different approach is adopted by [6] with the NOT algorithm. By choosing the narrowest interval for which the value of the chosen contrast function exceeds a pre-defined threshold, there is exactly one change-point in each interval with high probability. The ID method in [2] achieves, first, isolation and then detection of each change-point through an idea based on expanding intervals in a sequential way, starting from both the beginning and the end of the data sequence, in an interchangeable way. The isolation of each change-point maximizes the detection power. [32] proposes SeedBS which uses a deterministic construction of search intervals which can be pre-computed while [18], [20] and [33] consider moving sum statistics. [24] achieves a multiscale decomposition of the data with respect to an adaptively chosen unbalanced Haar wavelet basis, using a transformation of the data called TGUH transform and [10], through a ‘bottom-up’ approach, explores reverse segmentation, which involves the creation of a ‘solution path’ by deleting the change-point with the smallest CUSUM value in the segment determined by its closest left and right neighbors, in order to obtain a hierarchy of nested models. A method that does not fall in any of the two categories described above is FDRSeg, proposed by [34], which controls the false discovery rate in the sense that the number of falsely detected change-points is bounded linearly by the number of true jumps.
@@
Methods for detecting more general structural changes can again be split into the same two categories. Focusing on optimization-based methods, [5] considers the estimation of linear models based on the least squares principle. [31] introduces a trend filtering (TF) approach to produce piecewise-linear trend estimates using an $\ell_{1}$ penalty. [21] presents a locally dynamic approach, called CPOP, that finds the best continuous piecewise-linear fit to the data using penalized least squares and an $\ell_{0}$ penalty. [22] proposes the MARS method for flexible regression using splines functions with the degree and knot locations determined by the data. Two methods for optimizing the knot locations and smoothing parameters for least-squares or penalized splines are introduced by [46]. [50] introduces the frequentist information criterion for change-point detection which can detect changes in the mean, slope, standard deviation or serial correlation structure of a signal whose noise can be modeled by Gaussian, Wiener, or Ornstein-Uhlenbeck processes. [45] introduces PLANT, a bottom-up type algorithm that finds the points at which there is potential variation of the slope using a likelihood-based approach, and then recursively merges the adjacent segments. Considering algorithms using a suitably chosen contrast function, the algorithm NOT can be applied to both piecewise-linear and quadratic signals, while ID can be applied to piecewise-linear and can be extended to piecewise-quadratic signals. Finally, [35] extends TGUH to piecewise-linear signals and [29] proposes a moving sum methodology.
@@
In this project, we propose a data-adaptive change-point detection method, called Data Adaptive ISolation (labelled DAIS), that attempts to isolate the change-points before detection, by also taking into account the potential true locations of the change-points. This means that the algorithm does not start from a random interval, or the beginning/end of the signal, as most algorithms in the literature, but instead starts checking around a point that there is reason to believe it could be around a true change-point. The idea behind the data-adaptive nature of the algorithm, and the belief that the algorithm will start close to a change-point, arise from the fact that, for example, in the case that the unobserved true signal $f_{t}$ in (1) is piecewise-constant, taking pairwise differences between consecutive time steps will reveal a constant signal with the value 0, with spikes where the change-points occur. This means that the largest spike, in absolute value, occurs at the location of the change-point with the largest jump in the sequence, therefore the most prominent change-point. Similarly, in the case that $f_{t}$ is continuous and piecewise-linear, differencing the signal twice will again reveal a signal with spikes near to the true locations of the change-points. We try to use this fact in the observed signal, $X_{t}$. In the algorithm, we identify the location of this spike, which we will be referring to as the ‘largest difference’ for the rest of the paper, and test around it for possible change-points. A discussion about the connection between the isolation aspect of the algorithm and the location of the largest difference can be found in Section 2.3, while Section 3.3 complements this discussion with theoretical results in the case that $\epsilon_{t}$ is Gaussian. We want to emphasise that since the largest difference is only used as the starting point of the search for change-points in the data sequence, the true change-points are detected even when the largest difference is far from them. The novelty of our work lies in the data-adaptive nature of the algorithm as just described.
[…]
The data-adaptive nature of DAIS is what differentiates it from its competitors. A key methodological difference with other algorithms is the way the intervals are being checked. WBS and NOT check to the left and right of the interval in which detection has already occurred, using pre-drawn intervals, while WBS2 draws new intervals. Essentially, for both algorithms, the data sequence is split into three parts, one being the interval including the change-point that has just been detected, one to its left and one to its right. After detection, ID re-starts on the left or the right of the interval where detection occurred, splitting the data sequence in two parts. DAIS, similarly to WBS and WBS2, checks intervals both to the left and right of the interval in which detection occurred. However, every time it re-starts, expanding intervals around the location of the largest difference, in absolute value, are created and thus detection of the change-point in an interval in which it is isolated occurs with high probability. The data-adaptivity also gives an advantage in terms of computational complexity, as is explained in Section 4.1.
@@
The paper is organized as follows. Section 2 provides details on the methodology of DAIS. We present a simple example followed by a detailed explanation of the algorithm and, in Section 2.3, we discuss the connection between the location of the largest difference calculated at each step of the algorithm and the guarantee that the change-point can be detected in an interval in which it is isolated. In Section 3, we provide theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations as well as some theoretical results concerning the discussion of Section 2.3. Section 4 includes an explanation of how some parameters of DAIS are selected and a note on computational complexity, which is also compared to that of competitors. In Section 5, we provide results on simulated data and a comparison to state-of-the-art competitors. Section 6 discusses extensions to the algorithm in more complicated signal structures, which include temporal correlation, relaxing the Gaussianity assumption on the noise term $\epsilon_{t}$, and multivariate data sequences. In Section 7, two examples using real world data are presented. The first one is about crime data reported daily and the second one involves weekly data on the Euro to British pound exchange rate. Section 8 concludes the paper with a summary of the most important findings. In Appendix A, we provide the signals used in the simulation study of Section 5 and in Appendix B further simulation results can be found. Finally, the proofs of Theorems 1 and 2 and the outline of the proof of Theorem 3 can be found in Appendix C, D and E, respectively.
@@ 2.1 Simple example
We begin by presenting a toy example of how DAIS works in practice in the case of $f_{t}$ in (1) being a piecewise-constant signal. We have a sequence with length $T=100$, standard deviation $\sigma=1$, and a change-point at location $r=65$ with a jump of magnitude equal to $1.5$. The data are presented in the left plot of Figure 1 with the true underlying signal plotted in red. We first calculate the absolute difference of consecutive observations, $X_{t}$ and $X_{t+1}$ for $t=1,2,\dots,T-1$. As can be seen by the plot on the right of Figure 1, in this example the largest difference occurs at the location of the change-point ($d_{1,100}=65$), which means that $\textrm{argmax}_{t=1,2,\ldots,99}\lvert X_{t+1}-X_{t}\rvert=65$. As shown in the left plot of Figure 2, DAIS creates right- and left-expanding intervals around this point, where, for the sake of presentation of this example, the expansion parameter is chosen to be $\lambda_{T}=10$. The intervals in the order in which they are checked are $[65,74]$, $[55,74]$ and $[55,84]$, where detection occurs. Note here that the algorithm does not detect the change-point in the first two intervals being checked, most probably either because the magnitude of the jump compared to $\sigma$ is small, or because the number of observations used is not sufficiently large, or a combination of the two. The change-point is identified at location 65 and then the algorithm restarts in the intervals $[1,55]$ and $[84,100]$. As these two intervals contain no change-points in our example, the largest differences are just at points where the absolute value of the difference between the noise of two consecutive observations is large. In this case, the location of the largest difference is arbitrary and only acts as the starting point for searching the whole interval for change-points to confirm that there is none. The largest differences are now detected at points $d_{1,55}=46$ and $d_{84,100}=85$ in the two intervals, respectively. The algorithm starts checking the interval $[1,55]$, with the expansions being performed in a similar way as to the smaller interval $[84,100]$, on which we focus for simplicity. For the interval $[84,100]$ the left- and right-expanding subintervals checked are: $[85,94]$, $[84,94]$, and $[84,100]$, as shown in the right plot of Figure 2. Since no change-point is detected in either of the intervals, the algorithm terminates.
@@
Figure 1: Left plot: The data (in black) and the underlying signal (in red) used for the toy example. The change-point is at location 65. Right plot: The absolute values of consecutive pairwise differences. The maximum, with location $d_{1,100}=65$, is marked with a red circle.
@@
Figure 2: Example with one change-point at $r=65$ with jump 1.5. Left plot: The largest difference is at $d_{1,100}=65$ and the change-point is detected when the third expanding interval is checked. After detection the next intervals that are checked are $[1,55]$ and $[84,100]$. Right plot: The left- and right-expanding intervals around the location of the largest difference, $d_{84,100}=85$, when checking $[84,100]$. No change-point is detected.
We emphasise that the differences are only used as a way to start the algorithm in an area close to the change-point. In cases with high variance, $\sigma^{2}$, of the noise, the largest difference might not be around the change-point but under the assumptions specified in Section 3, the change-points will still be detected with high probability. The location of the largest difference that guarantees detection of the change-point in an interval in which it is isolated is discussed in Section 2.3, and further results are provided in Section 3.3 for the case that $\epsilon_{t}$ in (1) is Gaussian.